OpenStudy (anonymous):

Step 1: Pick a friend or family member to be the character of your word problem. This friend or family member may do one of the following: Drive a boat Drive a jet ski Step 2: Select a current speed of the water in mph. Step 3: Select the number of hours (be reasonable please) that your friend or family member drove the boat or jets ski against the current speed you chose in step 2. Step 4: Select the number of hours that your friend or family member made the same trip with the current (this should be a smaller number, as your friend or family member will be traveling wi

OpenStudy (anonymous):

Step 5: Write out the word problem you created and calculate how fast your friend or family member was traveling in still water. Round your answer to the nearest mph.

OpenStudy (anonymous):

good lord

OpenStudy (anonymous):

I know but i am not good with coming up with word problems

OpenStudy (anonymous):

lets make it easier make the current speed 10 mph george boats against the current for 3 hours and with the current for two hours

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

now we want the speed of the boat in still water. lets call this rate R, so rate with current is R+10 and rate against current is R- 10 since distance equals rate times time, and we know the distance there and back are the same, we can write \[D=(r+10)\times 2=(r-10)\times 3\] and solve for rate

OpenStudy (anonymous):

\[2r+20=3r-30\] \[20=r-30\] \[r=50\] nice and easy

OpenStudy (anonymous):

that's the answer ?

OpenStudy (anonymous):

answer is rate in still water is 50 mph so rate with the current was 60 mph and against the current was 40 mph

OpenStudy (anonymous):

we can even find the distance traveled if you like since \[D=RT\] we should have \[D=2\times 60=120\] and also \[D=3\times 40 =120\] so we have checked the answer and it works

OpenStudy (anonymous):

Thank You so much!